OpenStudy (anonymous):

If f(x) = 2x + 3 and g(x) = 4x – 5, what is f(g(x))

OpenStudy (agent47):

f(g(x))=f(4x-5)=2(4x-5)+3

OpenStudy (agent47):

=8x-7

OpenStudy (anonymous):

thanks